Forty members of a parent's organization are making candles to raise money. One member drops out and the rest have to make 3 more candles each to make up. Each member makes the same number of candles. How many candles do they make altogether?
Answer by ankor@dixie-net.com(22740)

Forty members of a parent's organization are making candles to raise money.
One member drops out and the rest have to make 3 more candles each to make up.
Each member makes the same number of candles.
How many candles do they make altogether?
:
Let c = original number of candles that each parent planned make
then
(c+3) = actual number of candles that were made
then
40c = the total number of candles required
:
39(c+3) = 40c
39c + 117 = 40c
117 = 40c - 39c
117 = c
:
Each parent planned to make 117 candles, 40 * 117 = 4680 altogether
:
Check this by finding how many 39 make; 120 * 39 = 4680